a) Find p so that the curve y = x 3 + 2 x 2 + p x + 3 has one horizontal tangent line only.

b) Find the value of x where this tangent line occurs.

Grade:12

1 Answers

SAGAR SINGH - IIT DELHI
13 years ago

Dear nitin,

  • a) The horizontal tangent line has a slope equal to zero. We first need to find the first derivative of y and equate to 0.

    y ' = 3 x 2 + 4 x + p = 0

  • For the above quadratic function to have one solution only, it needs to have its discriminant D equal to zero.

    D = 16 - 12 p = 0

  • Solve the above for p to find

    p = 4 / 3

  • b) To find the value of x where the horizontal tangent line occurs, we need to solve the above quadratic equation 3 x 2 + 4 x + p = 0 for x. It had the discriminant equal to zero, hence

    x = - 4 / 6 = - 2 / 3

We are all IITians and here to help you in your IIT JEE preparation.

All the best.

 If you like this answer please approve it....

win exciting gifts by answering the questions on Discussion Forum

 

Sagar Singh

B.Tech IIT Delhi